# RoPINN: Region Optimized Physics-Informed Neural Networks

### v2

@@ 1 Introduction
Although deep models have been proven with universal approximation capability, the actual optimization process of PINNs still faces thorny challenges [8, 24, 35]. As a basic topic of PINNs, the optimization problem has been widely explored from various aspects [17, 44]. Previous methods attempt to mitigate this problem by using novel architectures to enhance model capacity [58, 3, 50], reweighting multiple loss functions for more balanced convergence [47], resampling data to improve important areas [51] or developing new optimizers to tackle the rough loss landscape [55, 37], etc.
@@ A PDE with equation constraints, initial (ICs) and boundary conditions (BCs) can be formalized as
As a classical tool in traditional PDE solvers, the variational formulation is widely used to reduce the smoothness requirements of the approximation solution [42]. Concretely, the target PDEs are multiplied with a set of predefined test functions $\{v_{1},\cdots,v_{M}\}$ and then the PDE equation term of the loss function is transformed as follows [18, 19, 20]:
@@ 3.1 Region Optimization
For clarity, we record the point optimization loss defined in Eq. (2) at $\boldsymbol{x}$ as $\mathcal{L}(u_{\theta},\boldsymbol{x})$, where $\boldsymbol{x}\in\Omega\cup\partial\Omega$ denotes the point selected from inner domain, initial state or boundaries. We adopt $\mathcal{S}$ to denote the finite set of selected points. Then Eq. (2) can be simplified as $\mathcal{L}(u_{\theta},\mathcal{S})=\frac{1}{|\mathcal{S}|}\sum_{\boldsymbol{x}\in\mathcal{S}}\mathcal{L}(u_{\theta},\boldsymbol{x})$.
@@
Suppose that the loss function $\mathcal{L}$ is $L$-Lipschitz-$\beta$-smooth for $\theta$. If we run stochastic gradient method with step size $\alpha_{t}$ at the $t$-th step for $T$ iterations, we have that:
@@
Suppose that the point optimization loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth for $\theta$. If we run stochastic gradient method with step size $\alpha_{t}$ for $T$ iterations based on region optimization loss $\mathcal{L}_{r}^{\mathrm{region}}$ in Eq. (5), the generalization error in expectation satisfies:
@@
Based on the Lipschitz assumption, $\mathcal{E}_{\mathrm{gen}}$ can be bounded by times the expectation of distance between parameter $\theta$ optimized from different training sets. Region optimization paradigm will bring a more “consistent” gradient optimization direction than point optimization at each iteration, thereby benefitting the generalization property. See Appendix A.3 for complete proof. ∎
From Theorems 3.3 and 3.5, we can observe that region optimization can reduce the generalization error $\mathcal{E}_{\mathrm{gen}}$. Furthermore, the region optimization theorem also provides a more general theoretical framework. For example, the conventional point optimization is equivalent to the case of $\Omega_{r}=0$, where only one single point is selected for each region. For another extreme case, enlarging the region size to the whole domain (i.e. $\Omega_{r}=\Omega$), Eq. (5) is equivalent to directly optimizing the loss defined on $\Omega$, where the generalization error will be reduced to zero. Unfortunately, this ideal situation cannot be satisfied in practice, since Eq. (5) requires the precise calculation of the integral of the whole domain. More discussions about practical implementation are deferred to the next section.
@@
Corollary 3.6 implies that the integral on the input domain in region optimization can help training PINNs with high-order constraints, which is valuable for high-order PDEs, such as wave equations. In contrast, this valuable property cannot be achieved in the classic point optimization. See Example 3.7.
@@ 3.2 Practical Algorithm
Derived from our theoretical insights of region optimization, we implement RoPINN as a practical training algorithm. As elaborated in Algorithm 1, RoPINN involves the following two iterative steps: Monte Carlo approximation and trust region calibration, where the former can efficiently approximate the optimization objective and the latter is proposed to control the estimation error. Next, we will discuss the details and convergence properties of RoPINN. All proofs can be found in Appendix B.
@@
In addition to effectively approximating region optimization without adding sampling points, our proposed sampling-based strategy is also equivalent to a high-order loss function, especially for the first-order term, which is essential in practice [55]. Concretely, with Taylor expansion, we have that:
$\begin{split}\mathbb{E}_{\boldsymbol{\xi}\sim{U(\Omega_{r})}}\big{(}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})\big{)}=\mathbb{E}_{\boldsymbol{\xi}{\sim U(\Omega_{r})}}\left(\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x})+{\nabla_{\theta}(\boldsymbol{\xi}^{\sf T}\mathcal{L}_{1}}(u_{\theta},\boldsymbol{x}))+\mathcal{O}({\|\boldsymbol{\xi}\|^{2}})\right),\end{split}$ (9)
@@
  Input: number of iterations $T$, number of past iterations $T_{0}$ retained to estimate the trust region, default region size $r$, initial PINN parameters $\theta_{0}$ and trust region calibration value $\sigma_{0}=1$.
@@
Balance of Generalization and Optimization
Recall that in Theorem 3.5, we observe that a larger region size will benefit the generalization error, while Theorem 3.9 demonstrates that too large region size will also cause unstable training because it will result in excessive gradient estimation error of Monte Carlo sampling in our implementation. The above analysis reveals the underlying trade-off between generalization and optimization of PINN models, which is concretely formalized as follows.
@@
Based on the same assumption in Theorem 3.5 but optimize the model with the approximated region optimization loss $\mathcal{L}_{r}^{\mathrm{approx}}(u_{\theta},\boldsymbol{x})=\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi}),\boldsymbol{\xi}\sim U(\Omega_{r})$ for $T$ iterations, we further denote the upper bound of gradient estimation error as ${{\mathcal{E}_{r,\mathrm{grad}}}}=\max_{t\leq T}\|\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{approx}}-\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}\|$, then $\mathcal{E}_{\mathrm{gen}}$ satisfies:
@@
Compared to Theorem 3.5, considering the gradient estimation error will bring extra optimization discrepancy between different training sets. See Appendix B.4 for complete proof. ∎
Based on Theorem 3.12, we can find that classical point optimization ($\Omega_{r}=0$) and sampling points globally ($\Omega_{r}=\Omega$) discussed in Theorem 3.5 correspond to two extreme cases. The former makes ${\mathcal{E}_{r,\mathrm{grad}}}=0$ but cannot reduce the generalization error, while the latter one holds a large gradient estimation error. Thus, neither case can bring a perfect generalization. In contrast, the design for calibrating trust regions in RoPINN provides an adaptive strategy to better balance generalization and optimization, which can adjust the region size according to multi-iteration training stability.
@@ 4.2 Algorithm Analysis
To verify the effectiveness of our design in RoPINN, we present ablations in Figure 5. It is observed that although we only sample one point, even fixed-size region optimization can also boost the performance of PINNs in most cases, demonstrating the effectiveness of introducing “region” to PINN optimization. However, as illustrated in Theorem 3.9, the sampling process may also cause gradient estimation error, so the relative promotion is inconsistent and unstable among different PDEs and base models. With our proposed trust region calibration, we can obtain a more significant and consistent improvement, indicating that achieving a better balance between optimization and generalization (formalized in Theorem 3.12) performs an essential role in training PINN models.
@@ 5 Conclusion
This paper presents and analyzes a new PINN optimization paradigm: region optimization. Going beyond previous scatter-point optimization, we extend the optimization from selected points to their neighborhood regions. Based on this idea, RoPINN is implemented as a simple but effective training algorithm, where an effcient Monte Carlo approximation process is used along with a trust region calibration strategy to control the gradient estimation error caused by sampling, theoretically bringing a better balance of generalization and optimization. In addition to theoretical advantages, RoPINN can consistently boost the performance of various advanced PINN models without extra backpropagation or gradient calculation, demonstrating favorable efficiency, training stability and general capability.
@@
According to Lemma A.1, we attempt to bound the generalization error in expectation $\mathcal{E}_{\mathrm{gen}}$ by analyzing the error difference between two selected sample sets. Denote that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are two identical sample sets of size $|\mathcal{S}|$ except for one sample. Suppose that with the stochastic gradient method on these two sets, we can obtain two optimization trajectories $\{\theta_{t}\}_{t=1}^{T}$ and $\{\theta_{t}^{\prime}\}_{t=1}^{T}$ respectively. According to Assumption 3.2, we have the following inequality:
@@
This boundary is more tight than [13, 52], where the latter omits the $\frac{2L}{\beta(|\mathcal{S}|-1)}$ term in Eq. (27).
@@ D.5 Experiments with Advanced Quadrature Methods
In our implementation, RoPINN employs a simple Monte Carlo sampling to approximate integral. Obviously, we can adopt more advanced quadrature methods, such as Gaussian quadrature [16]. Thus, we also experiment with 2D space Gaussian quadrature, which requires square number points and the one-point-sampling situation will degenerate to the center value. As shown in Table 7, we can find that under our official setting (only sampling one point), Monte Carlo is better, while Gaussian quadrature is better in more points. Note that although Gaussian quadrature has the potential to achieve better performance, sampling more points may contradict our motivation to boost PINNs without extra backpropagation or gradient calculation. Thus, we choose the Monte Carlo method, which works better under high-efficiency settings.

### v3

@@ 1 Introduction
Although deep models have been proven to enjoy the universal approximation capability, the actual optimization process of PINNs still faces thorny challenges [8, 24, 35]. As a basic topic of PINNs, the optimization problem has been widely explored from various aspects [17, 44]. Previous methods attempt to mitigate this problem by using novel architectures to enhance model capacity [58, 3, 50], reweighting multiple loss functions for more balanced convergence [47], resampling data to improve important areas [51] or developing new optimizers to tackle the rough loss landscape [55, 37], etc.
@@ A PDE with equation constraints, initial (ICs) and boundary conditions (BCs) can be formalized as
As a classical tool in traditional PDE solvers, the variational formulation is widely used to reduce the smoothness requirements of the approximate solution [42]. Concretely, the target PDEs are multiplied with a set of predefined test functions $\{v_{1},\cdots,v_{M}\}$ and then the PDE equation term of the loss function is transformed as follows [18, 19, 20]:
@@ 3.1 Region Optimization
For clarity, we record the point optimization loss defined in Eq. (2) at $\boldsymbol{x}$ as $\mathcal{L}(u_{\theta},\boldsymbol{x})$, where $\boldsymbol{x}\in\Omega\cup{\Omega_{0}\cup}\partial\Omega$ denotes the point selected from inner domain, initial state or boundaries. We adopt $\mathcal{S}$ to denote the finite set of selected points. Then Eq. (2) can be simplified as $\mathcal{L}(u_{\theta},\mathcal{S})=\frac{1}{|\mathcal{S}|}\sum_{\boldsymbol{x}\in\mathcal{S}}\mathcal{L}(u_{\theta},\boldsymbol{x})$.
@@
Suppose that the loss function $\mathcal{L}$ is $L$-Lipschitz-$\beta$-smooth for $\theta$. If we run stochastic gradient descent with step size $\alpha_{t}$ at the $t$-th step for $T$ iterations, we have that:
@@
Suppose that the point optimization loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth for $\theta$. If we run stochastic gradient descent with step size $\alpha_{t}$ for $T$ iterations based on region optimization loss $\mathcal{L}_{r}^{\mathrm{region}}$ in Eq. (5), the generalization error in expectation satisfies:
@@
Based on the Lipschitz assumption, $\mathcal{E}_{\mathrm{gen}}$ can be bounded by a term relating to the expectation of distance between parameter $\theta$ optimized from different training sets. The region optimization paradigm brings a more “consistent” gradient optimization direction than point optimization at each iteration, thereby benefiting the generalization property. See Appendix A.3 for complete proof. ∎
From Theorems 3.3 and 3.5, we can observe that region optimization can reduce the generalization error $\mathcal{E}_{\mathrm{gen}}$. Furthermore, the region optimization theorem also provides a more general theoretical framework. For example, the conventional point optimization is equivalent to the case of $\Omega_{r}=0$, where only one single point is selected for each region. For another extreme case, enlarging the region size to the whole domain (i.e. $\Omega_{r}=\Omega$), Eq. (5) is equivalent to directly optimizing the loss defined on $\Omega$, where the generalization error will be reduced to zero. Unfortunately, this ideal situation cannot be satisfied in practice, since Eq. (5) requires precise calculation of the integral over the whole domain. More discussions of practical implementation are deferred to the next section.
@@
Corollary 3.6 implies that the integral on the input domain in region optimization can help training PINNs with high-order constraints, which is valuable for high-order PDEs, such as wave equations. In contrast, this valuable property cannot be achieved by the classic point optimization. See Example 3.7.
@@ 3.2 Practical Algorithm
Derived from our theoretical insights of region optimization, we implement RoPINN as a practical training algorithm. As elaborated in Algorithm 1, RoPINN involves the following two iterative steps: Monte Carlo approximation and trust region calibration, where the former can efficiently approximate the optimization objective and the latter can effectively control the estimation error. Next, we will discuss the details and convergence properties of RoPINN. All proofs can be found in Appendix B.
@@
In addition to efficiently approximating region optimization without adding sampling points, our proposed sampling-based strategy is also equivalent to a high-order loss function, especially for the first-order term, which is essential in practice [55]. Concretely, with Taylor expansion, we have that:
$\begin{split}\mathbb{E}_{\boldsymbol{\xi}\sim{U(\Omega_{r})}}\big{(}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi})\big{)}=\mathbb{E}_{\boldsymbol{\xi}{\sim U(\Omega_{r})}}\big{(}\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x})+{\nabla_{\theta}(\boldsymbol{\xi}^{\sf T}\mathcal{L}_{1}}(u_{\theta},\boldsymbol{x}))+\mathcal{O}({\|\boldsymbol{\xi}\|^{2}})\big{)},\end{split}$ (9)
@@
  Input: number of iterations $T$, number of past iterations $T_{0}$ retained to estimate the trust region, default region size $r$, trust region calibration value $\sigma_{0}=1$, and initial PINN parameters $\theta_{0}$.
@@
Balance between generalization and optimization
Recall that in Theorem 3.5, we observe that a larger region size will benefit the generalization error, while Theorem 3.9 demonstrates that too large region size will also cause unstable training because it will result in excessive gradient estimation error of Monte Carlo sampling in our implementation. The above analysis reveals the underlying trade-off between generalization and optimization of PINN models, which is formally stated below.
@@
Based on the same assumptions in Theorem 3.5 but optimizing the PINN model with the approximate region optimization loss $\mathcal{L}_{r}^{\mathrm{approx}}(u_{\theta},\boldsymbol{x})=\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x}+\boldsymbol{\xi}),\boldsymbol{\xi}\sim U(\Omega_{r})$ for $T$ iterations, we further denote the upper bound of gradient estimation error as ${{\mathcal{E}_{r,\mathrm{grad}}}}=\max_{t\leq T}\|\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{approx}}-\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}\|$, then $\mathcal{E}_{\mathrm{gen}}$ satisfies:
@@
In contrast to Theorem 3.5, here the gradient estimation error will bring extra optimization discrepancy between different training sets. See Appendix B.4 for complete proof. ∎
Based on Theorem 3.12, we have pinpointed that classical point optimization ($\Omega_{r}=0$) and sampling points globally ($\Omega_{r}=\Omega$) discussed in Theorem 3.5 correspond to two extreme cases. The former makes ${\mathcal{E}_{r,\mathrm{grad}}}=0$ but cannot reduce the generalization error, while the latter holds a large gradient estimation error. Thus, neither case can yield perfect generalization for PINNs . In contrast, the design for calibrating trust regions in RoPINN provides an adaptive strategy to better balance generalization and optimization, which can adjust the region size according to multi-iteration training stability.
@@ 4.2 Algorithm Analysis
To verify the effectiveness of our design in RoPINN, we present ablations in Figure 5. It is observed that although we only sample one point, even fixed-size region optimization can also boost the performance of PINNs in most cases, demonstrating the effectiveness of introducing “region” to PINN optimization. However, as illustrated in Theorem 3.9, the sampling process may also cause gradient estimation error, so the relative promotion is inconsistent and unstable among different PDEs and base models. With our proposed trust region calibration, we can obtain a more significant and consistent improvement, indicating that achieving a better balance between optimization and generalization (formally stated in Theorem 3.12) performs an essential role in training PINN models.
@@ 5 Conclusion
This paper presents and analyzes a new PINN optimization paradigm: region optimization. Going beyond previous scatter-point optimization, we extend the optimization from selected points to their neighborhood regions. Based on this idea, RoPINN is implemented as a simple but effective training algorithm, where an efficient Monte Carlo approximation process is used along with a trust region calibration strategy to control the gradient estimation error caused by sampling, theoretically manifesting a better balance of generalization and optimization. In addition to theoretical advantages, RoPINN can consistently boost the performance of various PINN models without extra backpropagation or gradient calculation, demonstrating favorable efficiency, training stability and general capability.
@@
According to Lemma A.1, we attempt to bound the generalization error in expectation $\mathcal{E}_{\mathrm{gen}}$ by analyzing the error difference between two selected sample sets. Denote $\mathcal{S}$ and $\mathcal{S}^{\prime}$ as two identical sample sets of size $|\mathcal{S}|$ except for one sample. Suppose that with the stochastic gradient method on these two sets, we can obtain two optimization trajectories $\{\theta_{t}\}_{t=1}^{T}$ and $\{\theta_{t}^{\prime}\}_{t=1}^{T}$ respectively. According to Assumption 3.2, we have the following inequality:
@@
This boundary is tighter than [13, 52], where the latter omits the $\frac{2L}{\beta(|\mathcal{S}|-1)}$ term in Eq. (27).
@@ D.5 Experiments with Advanced Quadrature Methods
In our implementation, RoPINN employs a simple Monte Carlo sampling to approximate integral. Obviously, we can adopt more advanced quadrature methods, such as Gaussian quadrature [16]. Thus, we also experiment with 2D space Gaussian quadrature, which requires square number points and the one-point-sampling situation will degenerate to the center value. As shown in Table 7, we can find that under our official setting (only sampling one point), Monte Carlo is better, while Gaussian quadrature is better in more points. Note that although Gaussian quadrature has the potential to achieve better performance, sampling more points may contradict our motivation of boosting PINNs without extra backpropagation or gradient calculation. Thus, we choose the Monte Carlo method, which works better under high-efficiency settings.